# Bias and Fairness in Large Language Models: A Survey

### v2

@@ 5.4.2 Discussion and Limitations
Post-processing mitigations do not assume access to a trainable model, which makes these appropriate techniques for black box models. That said, rewriting techniques are themselves prone to exhibiting bias. The determination of which outputs to rewrite is in itself a subjective and value-laden decision. Similar to potential harms with toxicity and sentiment classifiers (see Section 3.5.4), special care should be taken to ensure that certain social groups’ style of language is not disproportionately flagged and rewritten. The removal of protected attributes can also erase important contexts and produce less diverse outputs, itself a form of an exclusionary norm. Neural rewriters are also limited by the availability of parallel training corpora, which can restrict the dimensions of bias they are posed to address.
@@ 6 Open Problems & Challenges
In this section, we discuss open problems and highlight important challenges for future work.
@@ 6.1 Addressing Power Imbalances
Centering marginalized communities.
Technical solutions to societal injustices are incomplete, and framing technical mitigations as "fixes" to bias is problematic (Birhane, 2021; Byrum & Benjamin, 2022; Kalluri et al., 2020). Instead, technologists must critically engage with the historical, structural, and institutional power hierarchies that perpetuate harm and interrogate their own role in modulating those inequities. In particular, who holds power in the development and deployment of LLM systems, who is excluded, and how does technical solutionism preserve, enable, and strengthen inequality? Central to understanding the role of technical solutions – and to disrupting harmful power imbalances more broadly – is bringing marginalized communities into the forefront of LLM decision-making and system development, beginning with the acknowledgment and understanding of their lived experiences to reconstruct assumptions, values, motivations, and priorities. Researchers and practitioners should not merely react to bias in the systems they create, but instead design these technologies with the needs of vulnerable groups in mind from the start.
Developing participatory research designs.
@@
Reconstructing values and assumptions.
As we have established, bias and fairness are highly subjective and normative concepts situated in social, cultural, historical, political, and regional contexts. Therefore, there is no single set of values that bias and fairness research can assume, yet, as Green (2019) explains, the assumptions and values in scientific and computing research tend to reflect those of dominant groups. Instead of relying on vague notions of socially desirable behaviors of LLMs, researchers and practitioners can establish more rigorous theories of social change, grounded in relevant principles from fields like linguistics, sociology, and philosophy. These normative judgments should be made explicit and not assumed to be universal. One tangible direction of research in this vein is to expand bias and fairness considerations to contexts beyond the United States and Western ones often assumed by prior works, and for languages other than English. For example, several datasets rely on U.S. Department of Labor statistics to identify relevant dimensions for bias evaluation, which lacks generality to other regions of the world. Future work can expand perspectives to capture other sets of values and norms. Bhatt et al. (2022) and Malik et al. (2022) provide examples of such work for Indian society.
Expanding language resources.
Moving beyond the currently studied contexts will require additional language resources, including data for different languages and their dialects, as well as an understanding of various linguistic features and representations of bias. Curation of additional language resources should value inclusivity over convenience, and documentation should follow practices such as Bender & Friedman (2018) and Gebru et al. (2021). Furthermore, stakeholders must ensure that the process of collecting data itself does not contribute to further harm. Future work can examine frameworks for data collection pipelines that ensure communities maintain control over their own language resources and have a share in the benefits from the use of their data, following recommendations such as Jernite et al. (2022) and Walter & Suina (2019) to establish data governance and sovereignty practices.
@@ 6.2 Conceptualizing Fairness for NLP
Developing fairness desiderata.
We propose an initial set of fairness desiderata, but these notions can be refined and expanded. While works in machine learning classification have established extensive frameworks for quantifying bias and fairness, more work can be done to translate these notions and introduce new ones for NLP tasks, particularly for generated text, and for the unique set of representational harms that manifest in language. These definitions should stray away from abstract notions of fairness and instead be grounded in concrete injustices communicated and reinforced by language.
Rethinking social group definitions.
Delineating between social groups is often required to assess disparities, yet can simultaneously legitimize social constructions, reinforce power differentials, and enable systems of oppression (Hanna et al., 2020). Disaggregation offers a pathway to deconstruct socially constructed or overly general groupings, while maintaining the ability to perform disparity analysis. Disaggregated groups include intersectional ones, as well as more granular groupings of a population. Future work can leverage disaggregated analysis to develop improved evaluation metrics that more precisely specify who is harmed by an LLM and in what way, and more comprehensive mitigation techniques that take into account a broader set of social groups when targeting bias. In a similar vein, future work can more carefully consider how subgroups are constructed, as the definition of a social group can itself be exclusive. For example, Devinney et al. (2022) argue that modeling gender as binary and immutable erases the identities of trans, nonbinary, and intersex people. Bias and fairness research can expand its scope to groups and subgroups it has ignored or neglected. This includes supplementing linguistic resources like word lists that evaluation and mitigation rely on, and revising frameworks that require binary social groups. Another direction of research moves beyond observed attributes. Future work can interrogate techniques to measure bias for group identities that may not be directly observed, as well as the impact of proxies for social groups on bias.
Recognizing distinct social groups.
@@ 6.3 Refining Evaluation Principles
Establishing reporting standards.
Similar to model reporting practices established by Mitchell et al. (2019), we suggest that the evaluation of bias and fairness issues become standard additions to model documentation. That said, as we discuss throughout Section 3, several metrics are inconsistent with one another. For example, the selection of model hyperparameters or evaluation metric can lead to contradictory conclusions, creating confusing or misleading results, yet bias mitigation techniques often claim to successfully debias a model if any metric demonstrates a decrease in bias. Best practices for reporting bias and fairness evaluation remains an open problem. For instance, which or how many metrics should be reported? What additional information (e.g., evaluation dataset, model hyperparameters, etc.) should be required to contextualize the metric? How should specific harms be articulated? Han et al. (2023) provide a step in this direction, with an evaluation reporting checklist to characterize how test instances are aggregated by a bias metric. Orgad & Belinkov (2022) similarly outline best practices for selecting and stabilizing metrics. Works like these serve as a starting point for more robust reporting frameworks.
Considering the benefits and harms of universal benchmarks.
One possibility to standardize bias and fairness evaluation is to establish a universal benchmark. Universality can overcome comparability issues that arise from the vast array of bias evaluation metrics and datasets, enabling easier differentiation of bias mitigation techniques and their effectiveness. Despite this, universal benchmarks should be approached with caution. Universal benchmarks can obscure and decontextualize nuanced dimensions of harm, resulting in validity issues (Raji et al., 2021). In fact, universality may be completely at odds with the normative, subjective, and contextual nature of bias, and "universal" benchmarks may in actuality merely express the views of dominant groups and thus perpetuate further harm against marginalized groups. Future work can critically interrogate the role of a general evaluation framework, weighing the benefit of comparability with the risk of ineffectiveness.
Examining reliability and validity issues.
@@
Expanding evaluation possibilities.
This survey identifies and summarizes many different bias and fairness issues and their specific forms of harm that arise in LLMs. However, there are only a few such bias issues that are often explicitly evaluated, and for the ones that are, the set of evaluation techniques used for each type of bias remains narrow. For instance, most work leverages PerspectiveAPI for detecting toxicity despite the known flaws. Most work also relies on group fairness, with little emphasis towards individual or subgroup fairness. Additional metrics for each harm and notion of fairness should be developed and used.
@@ 6.4 Improving Mitigation Efforts
Developing hybrid techniques.
@@
Understanding mechanisms of bias within LLMs.
Some works like Jeoung & Diesner (2022) have examined how bias mitigation techniques change LLMs. For example, understanding that attention mechanisms play a key role in encoding bias informs attention-targeting mitigations such as Attanasio et al. (2022), Gaci et al. (2022), and Zayed et al. (2023a). Research into how and in which components (e.g., neurons, layers, attention heads, etc.) LLMs encode bias, and in what ways bias mitigations affect these, remains an understudied problem, with important implications for more targeted technical solutions.
@@ 6.5 Exploring Theoretical Limits
Establishing fairness guarantees.
@@
Analyzing performance-fairness trade-offs.
Bias mitigation techniques typically control a trade-off between performance and debiasing with a hyperparameter (e.g., regularization terms for in-training mitigations). Future work can better characterize this performance-fairness trade-off. For instance, Han et al. (2023) propose analysis of the Pareto frontiers for different hyperparameter values to understand the relationship between fairness and performance. We also refer back to our discussion of disaggregated analysis to carefully track what drives performance declines and whether performance changes are experienced by all social groups uniformly.

### v3

@@ 5.4.2 Discussion and Limitations
Post-processing mitigations do not assume access to a trainable model, which makes these appropriate techniques for black box models. That said, rewriting techniques are themselves prone to exhibiting bias. The determination of which outputs to rewrite is in itself a subjective and value-laden decision. Similar to potential harms with toxicity and sentiment classifiers (see Section 3.5.4), special care should be taken to ensure that certain social groups’ style of language is not disproportionately flagged and rewritten. The removal of protected attributes can also erase important contexts and produce less diverse outputs, itself a form of an exclusionary norm and erasure. Neural rewriters are also limited by the availability of parallel training corpora, which can restrict the dimensions of bias they are posed to address.
5.5 Recommendations
We synthesize findings and guidance from the literature to make the following recommendations. For more detailed discussion and limitations, see Sections 5.1.6, 5.2.5, 5.3.4, and 5.4.2.
  1. 1.
    Avoid flattening power imbalances. Data pre-processing techniques that rely on masking or replacing identity words may not capture the pertinent power dynamics that apply specifically and narrowly to certain social groups. If these techniques are deemed appropriate for the downstream application, ensure that the word lists are valid and complete representations of the social groups they intend to model.
  2. 2.
    Choose objective functions that align with fairness desiderata. Explicitly state the assumptions encoded in the choice of the loss or regularization function, or propose alternatives that are tailored to a specific fairness criterion. Consider cost-sensitive learning to increase the weight of minority classes in the training data.
  3. 3.
    Balance bias mitigation with output diversity. Ensure that minoritized voices are not filtered out due to modified decoding strategies. Rigorously validate that any heuristic intended to detect toxic or harmful tokens does not further marginalize social groups or their linguistic dialects and usages.
  4. 4.
    Preserve important contexts in output rewriting. Recognize the subjective and value-laden nature of determining which outputs to rewrite. Avoid flattening linguistic style and variation or erasing social group identities in post-processing.
@@ 6 Open Problems & Challenges
In this section, we discuss open problems and highlight challenges for future work.
@@ 6.1 Addressing Power Imbalances
Centering marginalized communities
Technical solutions to societal injustices are incomplete, and framing technical mitigations as "fixes" to bias is problematic (Birhane, 2021; Byrum and Benjamin, 2022; Kalluri et al., 2020). Instead, technologists must critically engage with the historical, structural, and institutional power hierarchies that perpetuate harm and interrogate their own role in modulating those inequities. In particular, who holds power in the development and deployment of LLM systems, who is excluded, and how does technical solutionism preserve, enable, and strengthen inequality? Central to understanding the role of technical solutions — and to disrupting harmful power imbalances more broadly — is bringing marginalized communities into the forefront of LLM decision-making and system development, beginning with the acknowledgment and understanding of their lived experiences to reconstruct assumptions, values, motivations, and priorities. Researchers and practitioners should not merely react to bias in the systems they create, but instead design these technologies with the needs of vulnerable groups in mind from the start (Grodzinsky, Miller, and Wolf, 2012).
Developing participatory research designs
@@
Shifting values and assumptions
As we have established, bias and fairness are highly subjective and normative concepts situated in social, cultural, historical, political, and regional contexts. Therefore, there is no single set of values that bias and fairness research can assume, yet, as Green (2019) explains, the assumptions and values in scientific and computing research tend to reflect those of dominant groups. Instead of relying on vague notions of socially desirable behaviors of LLMs, researchers and practitioners can establish more rigorous theories of social change, grounded in relevant principles from fields like linguistics, sociology, and philosophy. These normative judgments should be made explicit and not assumed to be universal. One tangible direction of research is to expand bias and fairness considerations to contexts beyond the United States and Western ones often assumed by prior works, and for languages other than English. For example, several datasets rely on U.S. Department of Labor statistics to identify relevant dimensions for bias evaluation, which lacks generality to other regions of the world. Future work can expand perspectives to capture other sets of values and norms. Bhatt et al. (2022) and Malik et al. (2022) provide examples of such work for Indian society.
Expanding language resources
Moving beyond the currently studied contexts will require additional language resources, including data for different languages and their dialects, as well as an understanding of various linguistic features and representations of bias. Curation of additional language resources should value inclusivity over convenience, and documentation should follow practices such as Bender and Friedman (2018) and Gebru et al. (2021). Furthermore, stakeholders must ensure that the process of collecting data itself does not contribute to further harms. As described by Jernite et al. (2022), this includes respecting the privacy and consent of the creators and subjects of data, providing people and communities with agency and control over their data, and sharing the benefits of data collection with the people and communities from whom the data originates. Future work can examine frameworks for data collection pipelines that ensure communities maintain control over their own language resources and have a share in the benefits from the use of their data, following recommendations such as Jernite et al. (2022) and Walter and Suina (2019) to establish data governance and sovereignty practices.
@@ 6.2 Conceptualizing Fairness for NLP
Developing fairness desiderata
We propose an initial set of fairness desiderata, but these notions can be refined and expanded. While works in machine learning classification have established extensive frameworks for quantifying bias and fairness, more work can be done to translate these notions and introduce new ones for NLP tasks, particularly for generated text, and for the unique set of representational harms that manifest in language. These definitions should stray away from abstract notions of fairness and instead be grounded in concrete injustices communicated and reinforced by language. For example, invariance (Definition 9), equal social group associations (Definition 10), and equal neutral associations (Definition 11) all represent abstract notion of consistency and uniformity in outcomes; it may be desirable, however, to go beyond sameness and instead ask how each social group and their corresponding histories and needs should be represented distinctly and uniquely to achieve equity and justice. The desiderata for promoting linguistic diversity to better represent the languages of minoritized communities in NLP systems, for instance, may differ from the desiderata for an NLP tool that assesses the quality of resumes in automated hiring systems. The desiderata and historical and structural context underpinning each definition should be made explicit.
Rethinking social group definitions
Delineating between social groups is often required to assess disparities, yet can simultaneously legitimize social constructions, reinforce power differentials, and enable systems of oppression (Hanna et al., 2020). Disaggregation offers a pathway to deconstruct socially constructed or overly general groupings, while maintaining the ability to perform disparity analysis within different contexts. Disaggregated groups include intersectional ones, as well as more granular groupings of a population. Future work can leverage disaggregated analysis to develop improved evaluation metrics that more precisely specify who is harmed by an LLM and in what way, and more comprehensive mitigation techniques that take into account a broader set of social groups when targeting bias. In a similar vein, future work can more carefully consider how subgroups are constructed, as the definition of a social group can itself be exclusive. For example, Devinney, Björklund, and Björklund (2022) argue that modeling gender as binary and immutable erases the identities of trans, nonbinary, and intersex people. Bias and fairness research can expand its scope to groups and subgroups it has ignored or neglected. This includes supplementing linguistic resources like word lists that evaluation and mitigation rely on, and revising frameworks that require binary social groups. Another direction of research moves beyond observed attributes. Future work can interrogate techniques to measure bias for group identities that may not be directly observed, as well as the impact of proxies for social groups on bias.
Recognizing distinct social groups
@@ 6.3 Refining Evaluation Principles
Establishing reporting standards
Similar to model reporting practices established by Mitchell et al. (2019), we suggest that the evaluation of bias and fairness issues become standard additions to model documentation. That said, as we discuss throughout Section 3, several metrics are inconsistent with one another. For example, the selection of model hyperparameters or evaluation metric can lead to contradictory conclusions, creating confusing or misleading results, yet bias mitigation techniques often claim to successfully debias a model if any metric demonstrates a decrease in bias. Best practices for reporting bias and fairness evaluation remain an open problem. For instance, which or how many metrics should be reported? What additional information (e.g., evaluation dataset, model hyperparameters, etc.) should be required to contextualize the metric? How should specific harms be articulated? Which contexts do evaluation datasets fail to represent and quantitative measures fail to capture? Han, Baldwin, and Cohn (2023) provide a step in this direction, with an evaluation reporting checklist to characterize how test instances are aggregated by a bias metric. Orgad and Belinkov (2022) similarly outline best practices for selecting and stabilizing metrics. Works like these serve as a starting point for more robust reporting frameworks.
Considering the benefits and harms of more comprehensive benchmarks
One possibility to standardize bias and fairness evaluation is to establish more comprehensive benchmarks to overcome comparability issues that arise from the vast array of bias evaluation metrics and datasets, enabling easier differentiation of bias mitigation techniques and their effectiveness. Despite this, benchmarks should be approached with caution and should not be conflated with notions of "universality." Benchmarks can obscure and decontextualize nuanced dimensions of harm, resulting in validity issues (Raji et al., 2021). In fact, overly general evaluation tools may be completely at odds with the normative, subjective, and contextual nature of bias, and "universal" benchmarks often express the perspectives of dominant groups in the name of objectivity and neutrality and thus perpetuate further harm against marginalized groups (Denton et al., 2020). Framing bias as something to be measured objectively ignores the assumptions made in the operationalization of the measurement tool (Jacobs and Wallach, 2021). It threatens to foster complacency when the benchmark is satisfied but the underlying power imbalance remains unaddressed. Future work can critically interrogate the role of a general evaluation framework, weighing the benefit of comparability with the risk of ineffectiveness.
Examining reliability and validity issues
@@
Expanding evaluation possibilities
This survey identifies and summarizes many different bias and fairness issues and their specific forms of harms that arise in LLMs. However, there are only a few such bias issues that are often explicitly evaluated, and for the ones that are, the set of evaluation techniques used for each type of bias remains narrow. For instance, most works leverage PerspectiveAPI for detecting toxicity despite the known flaws. Most works also rely on group fairness, with little emphasis towards individual or subgroup fairness. Additional metrics for each harm and notion of fairness should be developed and used.
@@ 6.4 Improving Mitigation Efforts
Developing hybrid techniques
@@
Understanding mechanisms of bias within LLMs
Some works like Jeoung and Diesner (2022) have examined how bias mitigation techniques change LLMs. For example, understanding that attention mechanisms play a key role in encoding bias informs attention-targeting mitigations such as Attanasio et al. (2022), Gaci et al. (2022), and Zayed et al. (2023a). Research into how and in which components (e.g., neurons, layers, attention heads, etc.) LLMs encode bias, and in what ways bias mitigations affect these, remains an understudied problem, with important implications for more targeted technical solutions.
@@ 6.5 Exploring Theoretical Limits
Establishing fairness guarantees
@@
Analyzing performance-fairness trade-offs
Bias mitigation techniques typically control a trade-off between performance and debiasing with a hyperparameter (e.g., regularization terms for in-training mitigations). Future work can better characterize this performance-fairness trade-off. For instance, Han, Baldwin, and Cohn (2023) propose analysis of the Pareto frontiers for different hyperparameter values to understand the relationship between fairness and performance. We also refer back to our discussion of disaggregated analysis in Section 6.1 to carefully track what drives performance declines and whether performance changes are experienced by all social groups uniformly. In this vein, we emphasize that achieving more fair outcomes should not be framed as an impediment to the standard, typically aggregated performance metrics like accuracy, but rather as a necessary criterion for building systems that do not further perpetuate harm.
7 Limitations
Technical solutions are incomplete without broader societal action against power hierarchies that diminish and dominate marginalized groups. In this vein, technical solutionism as an attitude overlooks and simplifies the broader histories and contexts that enable structural systems oppression, which can preserve, legitimate, and perpetuate the underlying roots of inequity and injustice, creating surface-level repairs that create an illusion of incremental progress but fail to interrogate or disrupt the broader systemic issues. This survey is limited in its alignment with a technical solutionist perspective, as opposed to a critical theoretical one. In particular, the taxonomies are organized according to their technical implementation details, instead of by their downstream usage contexts or harms. Though organization in this manner fails to question the broader and often tenuous assumptions in bias and fairness research more generally, we hope our organization can provide an understanding of the dominant narratives and themes in bias and fairness research for LLMs, enabling the identification of similarities between metrics, datasets, and mitigations with common underlying objectives and assumptions.
We have also focused narrowly on a few key points in the model development and deployment pipeline, particularly model training and evaluation. As Black et al. (2023) highlight, the decisions that researchers and practitioners can make in bias and fairness work are much more comprehensive. A more holistic approach includes problem formulation, data collection, and deployment and integration into real-world contexts.
Finally, this survey is limited in its focus on English language papers.